A salesperson has found that the probability of a sale on a single contact is approximately .03. If the salesperson contacts 100 prospects, what is the approximate probability of making at least one sale?


1
Expert's answer
2022-04-13T14:00:14-0400

"P(X \ge1)=1-P(0)=1-C^0_{100}(0.03)^0(0.97)^{100}=1-1x1x(0.97)^{100}=1-0.05=0.95"